OpenStudy (anonymous):

write a least degree polynomial whose zeros are 5i and -4

OpenStudy (precal):

3rd degree

OpenStudy (precal):

5i is a complex conjugate so therefore -5i is also a solution

OpenStudy (dumbcow):

\[x = \pm 5i\] \[x^2 = -25\] \[x^2 + 25 = 0\] ----------------- \[x = -4\] \[x +4 = 0\] Now we have factored form of polynomial \[(x^2 +25)(x+4)\] distribute to get standard form \[x^3 + 4x^2 +25x+100\]